OpenStudy (anonymous):

graph (x-1)^2/25-(y+2)^2/9=1

myininaya (myininaya):

do you realize this a hyperbola?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

hint \(\bf \cfrac{(x-{\color{brown}{ h}})^2}{{\color{blue}{ a}}^2}-\cfrac{(y-{\color{brown}{ k}})^2}{{\color{blue}{ b}}^2}=1 \\ \quad \\ \cfrac{(x-1)^2}{25}+\cfrac{(y+2)^2}{9}=1\implies \cfrac{(x-{\color{brown}{ 1}})^2}{{\color{blue}{ 5}}^2}-\cfrac{(y{\color{brown}{ +2}})^2}{{\color{blue}{ 3}}^2}=1\)

OpenStudy (jdoe0001):

get the center of it and use the traversal and conjugate axes to draw it bear in mind that the "positive fraction" variable is the axis where the hyperbola opens towards

OpenStudy (anonymous):

center is (1, -2) ?

OpenStudy (jdoe0001):

yeap center is 1,-2 traverse axis is "positive fraction" denominator.. .. so is 5 conjugate axis is 3

OpenStudy (anonymous):

so vertices are (6,-2) and (-4,-2) do i graph at these?

OpenStudy (jdoe0001):

well... yes.. you need the vertices to start the draw from and outwards

OpenStudy (anonymous):

okay thanks

OpenStudy (jdoe0001):

 notice the co-vertices are (-2,1) and (-2, -4)